# Summary of the NOTSOFAR-1 Challenge: Highlights and Learnings

### v1

@@ 6 Single-channel vs multi-channel accuracy
NOTSOFAR-1 challenge has confirmed the multi-channel systems’ ability to provide significantly more accurate transcription in challenging settings compared to single-channel systems. The winning multi-channel system by USTC [12] achieved relative improvement of 51% in terms of speaker attributed word error (tcpWER) and 46% in terms of speaker-agnostic word error rate (tcorcWER) over the winning single-channel system, also by USTC [12].
@@ 7.3 Inference techniques
  • Filtering out no speech segments to prevent hallucinations (NJU-AALab [14]).
@@ 8 Conclusions
This paper presented an overview of the NOTSOFAR-1 challenge, highlighting key results, insights and ideas for future improvement in DASR. The challenge captured the attention of the research community by introducing a dataset tailored not only to real-world scenarios but also to the pressing demands of business applications. With its combination of 280 recorded meetings across 30 distinct rooms and a simulated training dataset of 1,000 hours incorporating 15,000 real acoustic transfer functions (ATFs), the challenge set a new benchmark for realistic and diverse data in the field. This comprehensive dataset, paired with an open-source baseline system, attracted significant engagement from researchers, resulting in 11 submitted technical reports (9 from NOTSOFAR-1 and 2 from DASR tasks) that explored innovative solutions and demonstrated the potential for advancing DASR technologies. By bridging the gap between existing benchmarks and real-world applications, NOTSOFAR-1 established itself as a pivotal initiative, fostering collaboration and pushing the boundaries of speech processing research.

### v2

@@ 5 Single-channel systems
6 Dia-Sep-ASR vs CSS-ASR-Dia
In the previous sections we’ve concluded that Dia-Sep-ASR systems outperformed CSS-ASR-Dia systems in the challenge. In this section, we suggest several possible explanations to these observations. Without thorough ablation study of the systems, it is impossible to validate each of the offered explanations or estimate their relative impact, so we offer these potential explanations as hypotheses and encourage researchers to explore them in their future work.
While some observations apply to both SC and MC systems, the advantage of Dia-Sep-ASR systems was more evident in the MC track. The top three MC systems – USTC [14], STCON [12] and NTT [13] – are all Dia-Sep-ASR systems, demonstrating the effectiveness of this approach. The SC track was also won by a Dia-Sep-ASR system (USTC [14]); however, the absence of other high-performing Dia-Sep-ASR systems in this track (as STCON and NTT participated only in the DASR track and did not submit SC versions) makes the conclusion about Dia-Sep-ASR superiority less certain. Furthermore, the presence of USTC Enhanced Whisper [14] which likely contributed to USTC’s performance, further complicates direct comparisons.
6.1 Hypothesis 1: CSS is a more difficult task
Defined as a process of generating a set of overlap-free speech signals from a continuous audio stream consisting of multiple utterances spoken by different people [4], the CSS task inherently combines speaker diarization and speech separation. The baseline and submitted CSS-ASR-Dia systems tried to tackle both tasks using a single CSS model. In contrast, high-performing Dia-Sep-ASR systems employed complex multi-model diarization and separation subsystems. For example, USTC diarization subsystem alone incorporated two modified baseline CSS models – one repurposed for overlap detection and the other dedicated to speech separation for the sake of better speaker embeddings [14].
6.2 Hypothesis 2: longer context windows may have contributed to the superior performance of Dia-Sep-ASR systems
The baseline CSS model processed three-second speech segments, and submitted CSS-ASR-Dia systems maintained a similar window length. In contrast, top Dia-Sep-ASR systems incorporated components that processed much longer windows, potentially leveraging context more effectively. GSS, for instance, commonly operates on windows significantly longer than three seconds. Even STCON’s specialized “chunked” version of GSS, designed to accommodate frequent speaker movement, processed windows of about five seconds [12]. USTC’s T-F mask prediction network (JDS) and overlap detection component of their diarization subsystem operated on 12.8-second windows, while their cACGMM rectification component used a 120-second window. These differences suggest that CSS performance could improve with access to longer context windows.
6.3 Hypothesis 3: Dia-Sep-ASR systems may have been more effective in bridging gaps between simulated and real-world data
Despite significant efforts to make simulation data for model training as realistic as possible, some gaps between simulated and real data likely remain. We hypothesize that the submitted Dia-Sep-ASR systems may have been more effective in bridging these gaps by relying on GSS for speech separation, as it is less data-driven than neural network models. While neural networks generally benefit from better data utilization, simpler models like GSS may offer greater robustness when training and test data do not fully align. As discussed in Section 4.5, the submitted CSS models were trained solely on synthetic data. We believe that incorporating real-world data into CSS training could enhance its effectiveness and unlock the full potential of neural networks for speech separation.
@@ 7 Single-channel vs multi-channel accuracy
NOTSOFAR-1 challenge has confirmed the multi-channel systems’ ability to provide significantly more accurate transcription in challenging settings compared to single-channel systems. The winning multi-channel system by USTC [14] achieved relative improvement of 51% in terms of speaker attributed word error rate (tcpWER) and 46% in terms of speaker-agnostic word error rate (tcorcWER) over the winning single-channel system, also by USTC [14].
@@ 8.3 Inference techniques
  • Filtering out non-speech segments to prevent hallucinations (NJU-AALab [16]).
@@ 8.4 Re-scoring ASR output with language models (LM)
The winning system by USTC [14] didn’t use LM-based rescoring in their pipeline.
@@ 9 Conclusions
This paper presented an overview of the NOTSOFAR-1 challenge, highlighting key results, insights and ideas for future improvement in DASR. The challenge captured the attention of the research community by introducing a dataset tailored not only to real-world scenarios but also to the pressing demands of business applications. With its combination of 315 recorded meetings across 30 distinct rooms and a simulated training dataset of 1,000 hours incorporating 15,000 real acoustic transfer functions (ATFs), the challenge set a new benchmark for realistic and diverse data in the field. This comprehensive dataset, paired with an open-source baseline system, attracted significant engagement from researchers, resulting in 11 submitted technical reports (9 from NOTSOFAR-1 and 2 from DASR tasks) that explored innovative solutions and demonstrated the potential for advancing DASR technologies. By bridging the gap between existing benchmarks and real-world applications, NOTSOFAR-1 established itself as a pivotal initiative, fostering collaboration and pushing the boundaries of speech processing research.